f(x)=thesquarerootof2x-7dividedbythesquareroot1-x?whatisthedomain
Answer by venugopalramana(3286)

Y=SQRT(2X-7)/SQRT(1-X).....
DOMAIN IS THE VALUE OF X WHIH WILL MAKE THE FUNCTION MEANING FULL
HERE D.R.HAS SQRT(1-X)..WHICH IS ZERO AT X=1..........................I
SO THIS IS EXLUDED..
THEN SQRT(1-X)...WILL BE IMAGINARY IF 1-X IS NEGATIVE..SO 1-X>0..OR
1>X....OR.....X<1.....................................................II
NOW SQRT(2X-7)..ALSO NEEDS 2X-7>=0...OR....2X>=7...OR....X>=3.5...........III
NOW II AND III ARE CONTRADICTARY IF II IS TRUE THEN III IS FALSE AND VICEVERSA...HENCE DOMAIN OF X IS NULL SET..OR NO REAL VALUE